# ERDŐS–PÓSA PROPERTY OF CYCLES THAT ARE FAR APART

### v1

@@ 1 Introduction
We prove Theorem 1 for $f(k)\in\mathcal{O}(k^{18}\operatorname{polylog}k)$ and $g(d)=19d$. Note that the case $d=0$ of Theorem 1 with $g(0)=0$ is the original Erdős-Pósa Theorem (with suboptimal function $f$). The statement of Theorem 1 was also posed as a conjecture by Ahn et al. [1], who solved the cases $k=2$ and arbitrary $d$ as well as $d=1$ and arbitrary $k$.
@@ 3 The Proof
Let $r$ be a nonnegative integer, let $G$ be a graph, and let $C$ be a cycle in $G$. Then
@@
    $G$ has an $r$-unicyclic cycle $C^{\prime}$ with $V(C^{\prime})\subseteq B_{G}(V(C),2r)$; or
@@
    Both endpoints of $P$ are in $V(Q_{i})\setminus V(P_{i})$. In this case, we take $C_{i+1}$ to be the cycle in $Q_{i}\cup P$ and we take $Q_{i+1}:=Q_{i}\cap C_{i+1}$. This works because $P$ has two distinct endpoints in $Q_{i+1}$ and therefore $Q_{i+1}$ contains at least one edge, $Q_{i+1}\subseteq Q_{i}\subseteq C$, and $\operatorname{len}(P_{i+1})\leq 2r+1$ since $P_{i+1}=P$ in this case. Furthermore, $|Q_{i+1}|<|Q_{i}|$ because $C_{i+1}$ does not contain either endpoint of $Q_{i}$.
@@
This process eventually produces the desired cycle $C^{\prime}$ from the statement since, at each step in the process $|Q_{i}|$ decreases. ∎
@@
Let $d,r,k$ be integers such that $0\leq d\leq r$ and $k\geq 1$. Let $G$ be a graph and let $\mathcal{C}$ be a $2d$-packing of $d$-unicyclic cycles in $G$. Then either:
@@
      for each $\{C,C^{\prime}\}\in\binom{\mathcal{C}}{2}$, $B_{G}(V(C),r)\cap B_{G}(V(C^{\prime}),r)\subseteq Y$ and $Y$ contains at least one endpoint of each edge of $G$ between $B_{G}(V(C),r)$ and $B_{G}(V(C^{\prime}),r)$; and
@@
Let $d,r,k$ be integers such that $0\leq d\leq r$ and $k\geq 1$. Let $G$ be a graph and let $C$ be a $d$-unicyclic cycle in $G$. Let $U$ be a $C$-rooted spanning BFS-unicycle of $G_{r}:=G[B_{G}(V(C),r)]$. Then either:
@@
By Theorem 2, $G^{\prime}$ contains a set $\mathcal{D}$ of $k$ pairwise vertex-disjoint cycles. We now show that $\mathcal{D}$ is a $d$-packing of cycles in $G$. If $d=0$ then this is immediate, since the cycles in $\mathcal{D}$ are pairwise vertex-disjoint. Now suppose $d\geq 1$. Let $D$ and $D^{\prime}$ be two distinct cycles in $\mathcal{D}$. Let $v$ and $v^{\prime}$ be vertices of $D$ and $D^{\prime}$, respectively, such that $\operatorname{dist}_{G}(v,v^{\prime})=\operatorname{dist}_{G}(V(D),V(D^{\prime}))$. We must show that $\operatorname{dist}_{G}(v,v^{\prime})>d$.
@@
Since $V(P)\cap V(C)=\{v\}$, $v^{\prime}\notin V(C)$. Therefore $v^{\prime}$ is an internal vertex of $V(P_{e^{\prime}})$ for some $e^{\prime}\in I\setminus J$ and $P_{e^{\prime}}\subseteq D^{\prime}$. Since $C$ is $d$-unicyclic in $G$, every path in $G[B_{G}(V(C),d)]$ from $V(C)$ to $v^{\prime}$ contains an endpoint of $P_{e^{\prime}}$. In particular, $P$ contains an endpoint of $P_{e^{\prime}}$. Since $V(P)\cap V(C)=\{v\}$, this implies that $v$ is an endpoint of $P_{e^{\prime}}$. Then $v$ is a vertex of $P_{e^{\prime}}\subseteq D^{\prime}$, which contradicts the fact that $D$ and $D^{\prime}$ are vertex disjoint.
@@
Let $r,d,k$ be integers with $0\leq 2d\leq r$ and $k\geq 1$. Let $G$ be a graph, let $C_{1}$ and $C_{2}$ be $d$-unicyclic cycles in $G$ such that $\operatorname{dist}_{G}(V(C_{1}),V(C_{2}))>2d$. For each $i\in[2]$ let $G_{i,r}:=G[B_{G}(V(C_{i}),r)]$, let $U_{i}$ be a $C_{i}$-rooted spanning BFS-unicycle of $G_{i,r}$ and let $Y_{i}\subseteq V(G)$ be such that each edge in $E(G_{i,r})\setminus E(U_{i})$ has an endpoint in $Y_{i}$. Then either:
@@
      $Y$ contains every vertex in $V(G_{1,r})\cap V(G_{2,r})$ and $Y$ contains at least one endpoint of every edge of $G$ between $V(G_{1,r})$ and $V(G_{2,r})$; and
@@
For each $i\in[2]$ and each $v\in V(U_{i})$, let $P_{i,v}$ be the shortest path from $v$ to $V(C_{i})$ in $U_{i}$. Let $Y_{0}:=B_{G}(V(C_{1}),r)\cap B_{G}(V(C_{2}),r)$. Let $Z$ be the set of edges of $G$ between $V(G_{1,r})$ and $V(G_{2,r})$. For each $e=uv\in Z$ with $u\in V(G_{1,r})$ and $v\in V(G_{2,r})$, define $P_{1,e}:=P_{1,u}$ and $P_{2,e}=P_{2,v}$. For each $x$ in $Y_{0}\cup Z$, define $Q_{x}:=P_{1,x}\cup\{x\}\cup P_{2,x}$ and define $P_{x}$ to be a shortest path from $V(C_{1})$ to $V(C_{2})$ in $Q_{x}$.
Let $x$ be an element of $Y_{0}\cup Z$. Since $\operatorname{dist}(V(C_{1}),V(C_{2}))>d\geq 0$, the cycles $C_{1}$ and $C_{2}$ are vertex disjoint. The path $P_{x}$ is a path between $V(C_{1})$ and $V(C_{2})$ of length at least $1$ and at most $\operatorname{len}(P_{1,x})+1+\operatorname{len}(P_{2,x})\leq 2r+1$. ($P_{x}$ does not necessarily contain the vertex/edge $x$.) We say that $x\in Y_{0}\cup Z$ is uninteresting if $P_{x}$ contains a $U_{i}$-descendant of some vertex in $Y_{i}$ for some $i\in[2]$, otherwise $x$ is interesting.
Let $H$ be an auxiliary graph whose vertex set is the set of interesting elements in $Y_{0}\cup Z$ and in which two distinct vertices $x$ and $x^{\prime}$ are adjacent if and only if $\operatorname{dist}_{G}(V(P_{x}),V(P_{x^{\prime}}))\leq d$. Let $I$ be a maximal independent set in $H$. We split the proof into two cases. If $|I|\geq\frac{1}{2}s(k)$ we will show that outcome a holds. Otherwise, we show that outcome b holds.
First consider the case in which $|I|\geq\tfrac{1}{2}s(k)$. Let $G^{\prime}:=C_{1}\cup C_{2}\cup\bigcup_{x\in I}P_{x}$. For distinct $x,x^{\prime}\in I$, $xx^{\prime}$ is not an edge of $H$, so $\operatorname{dist}_{G}(V(P_{x}),V(P_{x^{\prime}}))>d\geq 0$. This together with the fact that $C_{1}$ is vertex-disjoint from $C_{2}$ gives that $G^{\prime}$ contains only vertices of degree $2$ and degree $3$. Recall that for each $x\in Y_{0}\cup Z$, the path $P_{x}$ contains two different endpoints. Therefore, the degree-$3$ vertices of $G^{\prime}$ are exactly the endpoints of paths in $\{P_{x}:x\in I\}$. This implies that $G^{\prime}$ contains $2|I|\geq s(k)$ vertices of degree $3$. By Theorem 2, $G^{\prime}$ contains a set $\mathcal{D}$ of $k$ pairwise vertex-disjoint cycles.
We claim that $\mathcal{D}$ is a $d$-packing of cycles in $G$. If $d=0$, this is immediate, because the cycles in $\mathcal{D}$ are pairwise vertex-disjoint. Now suppose $d\geq 1$. Let $D$ and $D^{\prime}$ be two distinct cycles in $\mathcal{D}$ and let $v\in V(D)$ and $v^{\prime}\in V(D^{\prime})$ be such that $\operatorname{dist}_{G}(v,v^{\prime})=\operatorname{dist}_{G}(V(D),V(D^{\prime}))$. We must show that $\operatorname{dist}_{G}(v,v^{\prime})>d$.
@@
First suppose that $v^{\prime}\in V(Q^{\prime})$. Let $Q$ be the subpath of $Q^{\prime}$ from $w^{\prime}$ to $v^{\prime}$. Since $Q\subseteq Q^{\prime}\subseteq P_{x^{\prime}}$ contains no vertex in $Y_{1}$ and $V(Q^{\prime})\subseteq B_{G}(V(C_{1}),r)$, we have that $Q\subseteq U_{1}$. Since $U_{1}$ is a $C_{1}$-rooted spanning BFS-unicycle of $G_{1,r}$, $\operatorname{dist}_{G}(w^{\prime},v^{\prime})=\operatorname{dist}_{U_{1}}(w^{\prime},v^{\prime})=|Q|$. Observe that $Q$ is a shortest path in $G$ between $V(C_{1})$ and $v^{\prime}$, and thus $|Q|\leq\operatorname{len}(P)\leq d$. It follows that $Q\subseteq G[B_{G}(V(C_{1}),d)]$. Since $C_{1}$ is $d$-unicyclic, all paths from $V(C_{1})$ to $v^{\prime}$ in $G[B_{G}(V(C_{1}),d)]$ contain $w^{\prime}$. In particular, $P\subseteq G[B_{G}(V(C_{1}),d)]$ is a path from $v\in V(C_{1})$ to $v^{\prime}$, so $w^{\prime}\in V(P)$. Since $V(P)\cap V(C_{1})=\{v\}$, this implies that $v=w^{\prime}$. However $v=w^{\prime}\in V(P_{y^{\prime}})\subseteq V(D^{\prime})$. This implies that $v$ is a vertex of both $D$ and $D^{\prime}$, which contradicts the fact that $D$ and $D^{\prime}$ are vertex-disjoint.
Next we consider the case when $v^{\prime}\notin V(Q^{\prime})$. Let $q$ be the last vertex of $Q^{\prime}$ and let $q^{\prime}$ be the neighbour of $q$ in $P_{y^{\prime}}$ that is not in $Q^{\prime}$. Then
$\operatorname{len}(P_{y^{\prime}})$ $\geq\operatorname{dist}_{G}(w^{\prime},q^{\prime})+\operatorname{dist}_{G}(q^{\prime},v^{\prime})+\operatorname{dist}_{G}(v^{\prime},V(C_{2}))$
@@
(The second inequality follows from $\operatorname{dist}_{G}(V(C_{1}),q^{\prime})\leq\operatorname{dist}_{G}(V(C_{1}),v^{\prime})+\operatorname{dist}_{G}(v^{\prime},q^{\prime})$ and $\operatorname{dist}_{G}(V(C_{1}),V(C_{2}))\leq\operatorname{dist}_{G}(V(C_{1}),v^{\prime})+\operatorname{dist}_{G}(v^{\prime},V(C_{2}))$.) This contradicts the fact that $\operatorname{len}(P_{x^{\prime}})\leq\operatorname{len}(P_{1,x^{\prime}})+\operatorname{len}(P_{2,x^{\prime}})+1\leq 2r+1$. This completes the proof (assuming $|I|\geq\tfrac{1}{2}s(k)$) that outcome a holds.
@@
be the set of degree-$3$ vertices in $G^{\prime}$. Note that $|X|=2\cdot|I|<s(k)$. We now define the set $Y_{Z}\subseteq V(G)$, define $Y:=Y_{0}\cup Y_{Z}$, and show that this choice of $X$ and $Y$ satisfies b.
Consider some element $x\in Y_{0}\cup Z$. If $x=y$ for some $y\in Y_{0}$ is a vertex and $y$ is a $U_{i}$-descendant of some vertex in $Y_{i}$ for some $i\in[2]$, then $y$ satisfies Item x. If $x=e$ is an edge in $Z$ and an endpoint of $e$ is a $U_{i}$-descendant of some vertex in $Y_{i}$ for some $i\in[2]$, then we add this endpoint of $e$ to the set $Y_{Z}$ and this newly added vertex satisfies Item x.
Therefore, we assume that this is not the case, which implies that $x$ is interesting, so $x\in V(H)$. Since $I$ is a maximal independent set in $H$, the set $I$ is also a dominating set. Therefore, either $x\in I$ or $x$ has a neighbor in $I$. In $G$, this translates into the statement: there exists $x^{\prime}\in I$ such that $\operatorname{dist}_{G}(V(P_{x}),V(P_{x^{\prime}}))\leq d$. (This includes the case $x=x^{\prime}$, which applies when $x\in I$.)
Let $x^{\prime}\in I$ be such that $\operatorname{dist}_{G}(V(P_{x^{\prime}}),V(P_{x}))\leq d$. Then $X$ contains the endpoints $u_{1}^{\prime}$ and $u_{2}^{\prime}$ of $P_{x^{\prime}}$. Then $V(P_{x^{\prime}})\subseteq V(P_{1,x^{\prime}}\cup P_{2,x^{\prime}})\subseteq B_{G}(\{u_{1}^{\prime},u_{2}^{\prime}\},r)$. Since $\operatorname{dist}_{G}(V(P_{x^{\prime}}),V(P_{x}))\leq d$, we have that $B_{G}(\{u_{1}^{\prime},u_{2}^{\prime}\},r+d)$ contains some vertex in $V(P_{x})\subseteq V(P_{1,x})\cup V(P_{2,x})$. If $x\in Y_{0}$ then $B_{G}(\{u_{1}^{\prime},u_{2}^{\prime}\},2r+d)$ contains $x$. If $x\in Z$ then $B_{G}(\{u_{1}^{\prime},u_{2}^{\prime}\},2r+d)$ contains an endpoint of $x$. We add this endpoint to the set $Y_{Z}$ and this newly added vertex satisfies Item y.
Recall that $Y:=Y_{0}\cup Y_{Z}$. We have just shown that $X$ and $Y$ satisfy ii. Furthermore, $Y_{0}=V(G_{1,r})\cap V(G_{2,r})$ and, for each $e\in Z$, $Y_{Z}$ contains an endpoint of $e$, so $Y$ satisfies i. Thus, the sets $X$ and $Y$ satisfy the conditions of outcome b. ∎
@@
If every cycle in $G$ contains a vertex in $B_{G}(\bigcup_{i\in[p]}V(C_{i}),4d)\cup B_{G}(\bigcup_{i\in[q]}V(D^{\prime}_{i}),4d)$ then the process ends. Otherwise, let $D$ be a cycle with no vertex in $B_{G}(\bigcup_{i\in[p]}V(C_{i}),4d)\cup B_{G}(\bigcup_{i\in[q]}V(D^{\prime}_{i}),4d)$. Apply Lemma 4 to $D$ to find a cycle $D^{\prime}$ such that either (a) $V(D^{\prime})\subseteq B_{G}(V(D),2d)$and $D^{\prime}$ is $d$-unicyclic in $G$; or (b) $V(D^{\prime})\subseteq B_{G}(V(D),3d)$and $|D^{\prime}|\leq 6d+2$. In case Item a, $V(D)\cap B_{G}(\bigcup_{i\in[p]}V(C_{i}),4d)=\emptyset$ and $V(D^{\prime})\subseteq B_{G}(V(D),2d)$, which implies that $V(D^{\prime})\cap B_{G}(\bigcup_{i\in[p]}V(C_{i}),2d)=\emptyset$. Therefore, $\{C_{1},\ldots,C_{p},D^{\prime}\}$ is a $2d$-packing of $d$-unicyclic cycles in $G$, which contradicts the choice of $\{C_{1},\ldots,C_{p}\}$.
@@
Let $\leavevmode\color[rgb]{0.76,0.13,0.28}X_{0}:=\{x_{1},\ldots,x_{q}\}$ where $x_{i}$ is an arbitrary vertex of $D^{\prime}_{i}$ for each $i\in[q]$. Since $|D_{i}^{\prime}|\leq 6d+2$, we have $V(D_{i}^{\prime})\subseteq B_{G}(x_{i},3d+1)$ for each $i\in[q]$. Therefore,

### v2

@@ 1 Introduction
We prove Theorem 1 for $f(k)\in\mathcal{O}(k^{18}\operatorname{polylog}k)$ and $g(d)=19d$. The statement of Theorem 1 was also posed as a conjecture by Ahn et al. [1], who solved the cases $k=2$ and arbitrary $d$ as well as $d=1$ and arbitrary $k$.
@@ 3 The Proof
Let $r$ be a nonnegative integer, let $G$ be a graph, and let $C$ be a cycle in $G$. Then either
@@
    $G$ has an $r$-unicyclic cycle $C^{\prime}$ with $V(C^{\prime})\subseteq B_{G}(V(C),2r)$, or
@@
    Both endpoints of $P$ are in $V(Q_{i})\setminus V(P_{i})$. In this case, we take $C_{i+1}$ to be the cycle in $Q_{i}\cup P$ and we take $Q_{i+1}:=Q_{i}\cap C_{i+1}$. This works because $P$ has two distinct endpoints in $Q_{i+1}$ and therefore $Q_{i+1}$ contains at least one edge, $Q_{i+1}\subseteq Q_{i}\subseteq C$, and $\operatorname{len}(P_{i+1})\leq 2r+1$ since $P_{i+1}=P$ in this case. Furthermore, $\operatorname{len}(Q_{i+1})<\operatorname{len}(Q_{i})$ because $C_{i+1}$ does not contain either endpoint of $Q_{i}$.
@@
This process eventually produces the desired cycle $C^{\prime}$ from the statement since, at each step in the process $\operatorname{len}(Q_{i})$ decreases. ∎
@@
Let $d,r,k$ be integers such that $1\leq d\leq r$ and $k\geq 1$. Let $G$ be a graph and let $\mathcal{C}$ be a $2d$-packing of $d$-unicyclic cycles in $G$. Then either:
@@
      for each $\{C,C^{\prime}\}\in\binom{\mathcal{C}}{2}$, $B_{G}(V(C),r)\cap B_{G}(V(C^{\prime}),r)\subseteq Y$; and
@@
Let $d,r,k$ be integers such that $1\leq d\leq r$ and $k\geq 1$. Let $G$ be a graph and let $C$ be a $d$-unicyclic cycle in $G$. Let $U$ be a $C$-rooted spanning BFS-unicycle of $G_{r}:=G[B_{G}(V(C),r)]$. Then either:
@@
By Theorem 2, $G^{\prime}$ contains a set $\mathcal{D}$ of $k$ pairwise vertex-disjoint cycles. We now show that $\mathcal{D}$ is a $d$-packing of cycles in $G$. Let $D$ and $D^{\prime}$ be two distinct cycles in $\mathcal{D}$. Let $v$ and $v^{\prime}$ be vertices of $D$ and $D^{\prime}$, respectively, such that $\operatorname{dist}_{G}(v,v^{\prime})=\operatorname{dist}_{G}(V(D),V(D^{\prime}))$. We must show that $\operatorname{dist}_{G}(v,v^{\prime})>d$.
@@
Since $V(P)\cap V(C)=\{v\}$, $v^{\prime}\notin V(C)$. Therefore $v^{\prime}$ is an internal vertex of $V(P_{e^{\prime}})$ for some $e^{\prime}\in I\setminus J$. Since every internal vertex of $P_{e^{\prime}}$ has degree $2$ in $G^{\prime}$, this implies that $P_{e^{\prime}}\subseteq D^{\prime}$. Since $C$ is $d$-unicyclic in $G$, every path in $G[B_{G}(V(C),d)]$ from $V(C)$ to $v^{\prime}$ contains an endpoint of $P_{e^{\prime}}$. In particular, $P$ contains an endpoint of $P_{e^{\prime}}$. Since $V(P)\cap V(C)=\{v\}$, this implies that $v$ is an endpoint of $P_{e^{\prime}}$. Then $v$ is a vertex of $P_{e^{\prime}}\subseteq D^{\prime}$, which contradicts the fact that $D$ and $D^{\prime}$ are vertex disjoint.
@@
Let $r,d,k$ be integers with $1\leq d\leq r$ and $k\geq 1$. Let $G$ be a graph, let $C_{1}$ and $C_{2}$ be $d$-unicyclic cycles in $G$ such that $\operatorname{dist}_{G}(V(C_{1}),V(C_{2}))>2d$. For each $i\in[2]$ let $G_{i,r}:=G[B_{G}(V(C_{i}),r)]$, let $U_{i}$ be a $C_{i}$-rooted spanning BFS-unicycle of $G_{i,r}$ and let $Y_{i}\subseteq V(G)$ be such that each edge in $E(G_{i,r})\setminus E(U_{i})$ has an endpoint in $Y_{i}$. Then either:
@@
      $Y$ contains every vertex in $V(G_{1,r})\cap V(G_{2,r})$; and
@@
For each $i\in[2]$ and each $v\in V(U_{i})$, let $P_{i,v}$ be the shortest path from $v$ to $V(C_{i})$ in $U_{i}$. Let $Y:=B_{G}(V(C_{1}),r)\cap B_{G}(V(C_{2}),r)$. For each $x$ in $Y$, define $Q_{x}:=P_{1,x}\cup\{x\}\cup P_{2,x}$ and define $P_{x}$ to be a shortest path from $V(C_{1})$ to $V(C_{2})$ in $Q_{x}$.
Let $x$ be an element of $Y$. Since $\operatorname{dist}(V(C_{1}),V(C_{2}))>d\geq 0$, the cycles $C_{1}$ and $C_{2}$ are vertex disjoint. The path $P_{x}$ is a non-trivial path between $V(C_{1})$ and $V(C_{2})$ of length at least $1$ and at most $\operatorname{len}(P_{1,x})+\operatorname{len}(P_{2,x})\leq 2r$. ($P_{x}$ does not necessarily contain the vertex $x$.) We say that $x\in Y$ is uninteresting if $P_{x}$ contains a $U_{i}$-descendant of some vertex in $Y_{i}$ for some $i\in[2]$, otherwise $x$ is interesting.
Let $H$ be an auxiliary graph whose vertex set is the set of interesting elements in $Y$ and in which two distinct vertices $x$ and $x^{\prime}$ are adjacent if and only if $\operatorname{dist}_{G}(V(P_{x}),V(P_{x^{\prime}}))\leq d$. Let $I$ be a maximal independent set in $H$. We split the proof into two cases. If $|I|\geq\frac{1}{2}s(k)$ we will show that outcome a holds. Otherwise, we show that outcome b holds.
First consider the case in which $|I|\geq\tfrac{1}{2}s(k)$. Let $G^{\prime}:=C_{1}\cup C_{2}\cup\bigcup_{x\in I}P_{x}$. For distinct $x,x^{\prime}\in I$, $xx^{\prime}$ is not an edge of $H$, so $\operatorname{dist}_{G}(V(P_{x}),V(P_{x^{\prime}}))>d\geq 1$. This together with the fact that $C_{1}$ is vertex-disjoint from $C_{2}$ gives that $G^{\prime}$ contains only vertices of degree $2$ and degree $3$. Recall that for each $x\in Y$, the path $P_{x}$ contains two different endpoints. Therefore, the degree-$3$ vertices of $G^{\prime}$ are exactly the endpoints of paths in $\{P_{x}:x\in I\}$. This implies that $G^{\prime}$ contains $2|I|\geq s(k)$ vertices of degree $3$. By Theorem 2, $G^{\prime}$ contains a set $\mathcal{D}$ of $k$ pairwise vertex-disjoint cycles.
We claim that $\mathcal{D}$ is a $d$-packing of cycles in $G$. Let $D$ and $D^{\prime}$ be two distinct cycles in $\mathcal{D}$ and let $v\in V(D)$ and $v^{\prime}\in V(D^{\prime})$ be such that $\operatorname{dist}_{G}(v,v^{\prime})=\operatorname{dist}_{G}(V(D),V(D^{\prime}))$. We must show that $\operatorname{dist}_{G}(v,v^{\prime})>d$.
@@
First suppose that $v^{\prime}\in V(Q^{\prime})$. Let $Q$ be the subpath of $Q^{\prime}$ from $w^{\prime}$ to $v^{\prime}$. Since $Q\subseteq Q^{\prime}\subseteq P_{x^{\prime}}$ contains no vertex in $Y_{1}$ and $V(Q^{\prime})\subseteq B_{G}(V(C_{1}),r)$, we have that $Q\subseteq U_{1}$. Since $U_{1}$ is a $C_{1}$-rooted spanning BFS-unicycle of $G_{1,r}$, $\operatorname{dist}_{G}(w^{\prime},v^{\prime})=\operatorname{dist}_{U_{1}}(w^{\prime},v^{\prime})=\operatorname{len}(Q)$. Observe that $Q$ is a shortest path in $G$ between $V(C_{1})$ and $v^{\prime}$, and thus $\operatorname{len}(Q)\leq\operatorname{len}(P)\leq d$. It follows that $Q\subseteq G[B_{G}(V(C_{1}),d)]$. Since $C_{1}$ is $d$-unicyclic, all paths from $V(C_{1})$ to $v^{\prime}$ in $G[B_{G}(V(C_{1}),d)]$ contain $w^{\prime}$. In particular, $P\subseteq G[B_{G}(V(C_{1}),d)]$ is a path from $v\in V(C_{1})$ to $v^{\prime}$, so $w^{\prime}\in V(P)$. Since $V(P)\cap V(C_{1})=\{v\}$, this implies that $v=w^{\prime}$. However $v=w^{\prime}\in V(P_{x^{\prime}})\subseteq V(D^{\prime})$. This implies that $v$ is a vertex of both $D$ and $D^{\prime}$, which contradicts the fact that $D$ and $D^{\prime}$ are vertex disjoint.
Next we consider the case when $v^{\prime}\notin V(Q^{\prime})$. Let $q$ be the last vertex of $Q^{\prime}$ and let $q^{\prime}$ be the neighbour of $q$ in $P_{x^{\prime}}$ that is not in $Q^{\prime}$. Then
$\operatorname{len}(P_{x^{\prime}})$ $\geq\operatorname{dist}_{G}(w^{\prime},q^{\prime})+\operatorname{dist}_{G}(q^{\prime},v^{\prime})+\operatorname{dist}_{G}(v^{\prime},V(C_{2}))$
@@
(The second inequality follows from $\operatorname{dist}_{G}(V(C_{1}),q^{\prime})\leq\operatorname{dist}_{G}(V(C_{1}),v^{\prime})+\operatorname{dist}_{G}(v^{\prime},q^{\prime})$ and $\operatorname{dist}_{G}(V(C_{1}),V(C_{2}))\leq\operatorname{dist}_{G}(V(C_{1}),v^{\prime})+\operatorname{dist}_{G}(v^{\prime},V(C_{2}))$.) This contradicts the fact that $\operatorname{len}(P_{x^{\prime}})\leq\operatorname{len}(P_{1,x^{\prime}})+\operatorname{len}(P_{2,x^{\prime}})\leq 2r$. This completes the proof (assuming $|I|\geq\tfrac{1}{2}s(k)$) that outcome a holds.
@@
be the set of degree-$3$ vertices in $G^{\prime}$. Note that $|X|=2\cdot|I|<s(k)$. We show that $X$ and $Y$ satisfy b.
Consider some vertex $y\in Y$. If $y$ is a $U_{i}$-descendant of some vertex in $Y_{i}$ for some $i\in[2]$, then $y$ satisfies Item x. Therefore, we assume that this is not the case, which implies that $y$ is interesting, so $y\in V(H)$. Since $I$ is a maximal independent set in $H$, the set $I$ is also a dominating set. Therefore, either $y\in I$ or $y$ has a neighbor in $I$. In $G$, this translates into the statement: there exists $y^{\prime}\in I$ such that $\operatorname{dist}_{G}(V(P_{y}),V(P_{y^{\prime}}))\leq d$. (This includes the case $y=y^{\prime}$, which applies when $y\in I$.)
Let $y^{\prime}\in I$ be such that $\operatorname{dist}_{G}(V(P_{y^{\prime}}),V(P_{y}))\leq d$. Then $X$ contains the endpoints $u_{1}^{\prime}$ and $u_{2}^{\prime}$ of $P_{y^{\prime}}$. Then $V(P_{y^{\prime}})\subseteq V(P_{1,y^{\prime}}\cup P_{2,y^{\prime}})\subseteq B_{G}(\{u_{1}^{\prime},u_{2}^{\prime}\},r)$. Since $\operatorname{dist}_{G}(V(P_{y^{\prime}}),V(P_{y}))\leq d$, we have that $B_{G}(\{u_{1}^{\prime},u_{2}^{\prime}\},r+d)$ contains some vertex in $V(P_{y})\subseteq V(P_{1,y})\cup V(P_{2,y})$, and thus $B_{G}(\{u_{1}^{\prime},u_{2}^{\prime}\},2r+d)$ contains $y$, as desired.
We have just shown that $X$ and $Y$ satisfy ii. Furthermore, $Y=V(G_{1,r})\cap V(G_{2,r})$. Thus, the sets $X$ and $Y$ satisfy the conditions of outcome b. ∎
@@
If every cycle in $G$ contains a vertex in $B_{G}(\bigcup_{i\in[p]}V(C_{i}),4d)\cup B_{G}(\bigcup_{i\in[q]}V(D^{\prime}_{i}),4d)$ then the process ends. Otherwise, let $D$ be a cycle with no vertex in $B_{G}(\bigcup_{i\in[p]}V(C_{i}),4d)\cup B_{G}(\bigcup_{i\in[q]}V(D^{\prime}_{i}),4d)$. Apply Lemma 4 to $D$ to find a cycle $D^{\prime}$ such that either (a) $V(D^{\prime})\subseteq B_{G}(V(D),2d)$and $D^{\prime}$ is $d$-unicyclic in $G$; or (b) $V(D^{\prime})\subseteq B_{G}(V(D),3d)$and $D^{\prime}$ is of length at most $26d+$ 2. In case Item a, $V(D)\cap B_{G}(\bigcup_{i\in[p]}V(C_{i}),4d)=\emptyset$ and $V(D^{\prime})\subseteq B_{G}(V(D),2d)$, which implies that $V(D^{\prime})\cap B_{G}(\bigcup_{i\in[p]}V(C_{i}),2d)=\emptyset$. Therefore, $\{C_{1},\ldots,C_{p},D^{\prime}\}$ is a $2d$-packing of $d$-unicyclic cycles in $G$, which contradicts the choice of $\{C_{1},\ldots,C_{p}\}$.
@@
Let $\leavevmode\color[rgb]{0.76,0.13,0.28}X_{0}:=\{x_{1},\ldots,x_{q}\}$ where $x_{i}$ is an arbitrary vertex of $D^{\prime}_{i}$ for each $i\in[q]$. Since $D_{i}^{\prime}$ has length at most $26d+$ 2, we have $V(D_{i}^{\prime})\subseteq B_{G}(x_{i},3d+1)$ for each $i\in[q]$. Therefore,